Ms anna receives a bill of N$45860 for 244 units of electricity and 315 units of water. last month, he was billed N$43922 for 310 units of electricity and 201 of water. it the unit price for these commodities did not change, how much would he play for 312 units of electricity and 294 unit of water?
Answer by stanbon(75887)

Ms anna receives a bill of N$45860 for 244 units of electricity and 315 units of water. last month, he was billed N$43922 for 310 units of electricity and 201 of water. it the unit price for these commodities did not change, how much would he pay for 312 units of electricity and 294 unit of water?
--------
Equations:
244e + 315w = 45860
310e + 201w = 43922
------
I used a matrix method on my TI-84 to get:
e = 95N per unit
w = 72N per unit
-----
Ans: 312*95 + 294*72 = N50808
===================================
Cheers,
Stan H.